# VD3D: Taming Large Video Diffusion Transformers for 3D Camera Control

### v1

@@ 3.2 Camera control for spatiotemporal transformers
This approach mirrors the technique used in CameraCtrl [19], a concurrent study focusing on camera control in U-Net-based video diffusion models. The motivation for using Plücker coordinates is that geometric manipulations in the Plücker space can be performed through simple arithmetics on the coordinates, which makes it easier for the network to use the positional information stored in such a disentangled representation.
@@ 5 Conclusion
Appendix A Quantitative evaluation
@@
A.1 Multi-view generation
Table 3: Multi-view generation. We evaluate all models using reference camera trajectories and single-view input images of the RealEstate10K test set. We compute reconstruction metrics based on the subsequent frames for the low-resolution and upsampled high-resolution generations.
Method Low-resolution High-resolution
PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$)
Base Model 14.74 0.320 0.334 13.23 0.459 0.572
MotionCtrl 15.07 0.348 0.308 13.42 0.467 0.560
MotionCtrl (frozen) 14.59 0.308 0.340 13.11 0.455 0.573
CameraCtrl 14.81 0.327 0.330 13.21 0.456 0.571
Ours 17.23 0.534 0.211 14.90 0.499 0.499
w/o Plucker 14.89 0.346 0.308 13.05 0.455 0.573
w/o ControlNet 14.66 0.313 0.340 13.10 0.450 0.573
w/o weight copy 16.96 0.509 0.220 14.75 0.495 0.504
@@ A.1 Multi-view generation
A.2 Quality metrics
Table 4: Quality metrics evaluation. We evaluate all models using text prompts from the RealEstate10K and MSR-VTT test sets respectively.
Method RealEstate10K MSR-VTT
FID ($\downarrow$) FVD ($\downarrow$) CLIPSIM ($\uparrow$) FID ($\downarrow$) FVD ($\downarrow$) CLIPSIM ($\uparrow$)
Base Model 8.22 160.37 0.2677 3.50 141.26 0.2774
MotionCtrl 1.50 52.30 0.2708 9.97 183.57 0.2677
MotionCtrl (frozen) 3.53 142.15 0.2772 8.19 165.48 0.2679
CameraCtrl 2.28 66.31 0.2730 8.47 181.90 0.2690
Ours 1.40 42.43 0.2807 7.80 165.18 0.2689
w/o Plucker 1.17 43.65 0.2715 9.84 152.91 0.2660
w/o ControlNet 3.66 137.06 0.2766 8.34 185.79 0.2674
w/o weight copy 1.38 42.00 0.2710 10.09 218.43 0.2647

### v2

@@ 3.2 Camera control for spatiotemporal transformers
This approach mirrors the technique used in recent 3D works [95, 12, 46] as well as CameraCtrl [29], a concurrent study focusing on camera control in U-Net-based video diffusion models. The motivation for using Plücker coordinates is that geometric manipulations in the Plücker space can be performed through simple arithmetics on the coordinates, which makes it easier for the network to use the positional information stored in such a disentangled representation.
@@ 5 Conclusion
Appendix A Related Work
3D generation.
Limited to single category trainings, early work on 3D generation extends GANs into 3D using a neural renderer as an inductive bias [18, 10, 76, 89, 3], Towards more diverse and flexible generation, CLIP-based supervision [84] enabled text-based generation and editing of 3D assets [14, 42, 88, 43, 25, 113]. With recent advances in diffusion models, Score Distillation Sampling (SDS) replaces CLIP supervision with diffusion model supervision  [80, 118, 61, 15, 60, 114, 58, 30, 138, 68, 144, 48, 53, 96] for higher quality generation. In order to improve the 3D structure of scenes, another line of work generates multiple views of a scene [62, 67, 91, 20, 66, 50, 111, 35]. Alternatively, other methods use iterative inpainting for 3D scene generation [36, 92]. Recent methods lift input images into 3D  [11, 102, 26, 69, 140, 103, 83, 70, 112, 99] using NeRF [74], 3D Gaussian Splatting [49], or Meshes in combination with diffusion models. Other recent methods [37, 56, 131, 130, 147, 28, 44, 126, 101, 104, 81, 100, 98] tackle fast feed-forward 3D generation, directly predicting a 3D generation from input images or text. Different from our approach, these methods can only synthesize static scenes.
Appendix B Quantitative evaluation
@@
B.1 Multi-view generation
@@
B.2 Quality metrics